OpenStudy (anonymous):

log2(5x^2-8x)=2 find x

OpenStudy (anonymous):

question...are you allowed to use a calculator or no?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

anybody know?

OpenStudy (anonymous):

apply base 2 exponential to both sides, it cancels the log 5x^2-8x=2^2

OpenStudy (anonymous):

so square both sides? i don't understand?

OpenStudy (anonymous):

hang on it looks like this right \[\log _{2}(5x^{2}-8x)=2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright, your log base is 2, to undo that you do this \[\large 2^{\log_{2}(5x^{2}-8x)}=2^2\]

OpenStudy (anonymous):

the left 2 and log2 cancel out and leaves 5x^2-8x=4

OpenStudy (anonymous):

x= 12/5?

OpenStudy (anonymous):

No, move the 4 over 5x^2-8x-4=0 it factors as (x-2)(5x+2) so solve x-2=0 and 5x+2=0

OpenStudy (anonymous):

x = 2 , x = -2/5

OpenStudy (anonymous):

Right